OpenStudy (anonymous):

WILL GIVE MEDAL & FAN! If a plant manufacturing carpet backing has a cylindrical tank that is 3 meters high and is 6 meters in diameter. What is the volume of the tank? Assume 3.14 for pi. A) 28.3 m3 B) 37.7 m3 C) 56.5 m3 D) 84.8 m3

OpenStudy (anonymous):

What do you think it is?

OpenStudy (tfht):

the equation to find the volume of a cylinder is \[\pi*r ^{2}*h\] so if you plug in the correct values, you should have an equation of \[\pi*3 ^{2}*3\] which is equal to \[\pi*27\] which equals approximately 84.823 meters^3